A couple wants to have 2 boys and 2 girls. They will continuous to have children until they have reached the goal. Assuming that the probability that they have 54% chance to have a boy and 46% chance to have a girl in every pregnancy. No twin or more are possible. Assuming that their goal is possible and has been fulfilled today! a. What is the probability that they will have more than 4 children?
Answer #1

Ans:

P(more than 4 children)=1-P(less than equal to 4)

(less than 4 is not possible,so there is only one option for meeting goal,they have 4 children i.e. 2 boys and 2 girls))

=1-0.54*0.54*0.46*0.46

=1-0.0617

   =0.9383